A rectangular loop of wire with sides H-26 cm and w-79 cm carries current . 0.344 A. An infinite straight wire, located a distance L-20 cm from segment ad of the loop as shown, carries current 1 -0.575 A in the positive y-direction. I2 a W I 1) What is Fad,x, the x component of the force exerted by the infinite wire on segment ad of the loop? N Submit 2) What is Fbc the x-component of the force exerted by the infinite wire on segment bc of the loop?. N Submit 3) What is Fnet.y, the y-component of the net force exerted by the infinite wire on the loop? Submit
media%2F085%2F085e187d-b011-4cdd-9f4e-cd
Answer #1

1.

Fadx = force on segment "ad" = (\mu/4\pi) (2I1 I2 H/L)

Fadx = (10-7) (2 x 0.575 x 0.344 x 0.26/0.20)

Fadx = 5.14 x 10-8 N

2.

Fbcx = force on segment "bc" = (\mu/4\pi) (2I1 I2 H/(L+ W))

Fbcx = (10-7) (2 x 0.575 x 0.344 x 0.26/(0.20 + 0.79))

Fbcx = 1.04 x 10-8 N

3)

FnetY = 0

since force on the upper section of wire is cancelled by the force on lower section of wire.

4)

the net force by the left wire on the loop is in positive x-direction toward right, hence the net force by the wire on right must be towards left and equal. hence the current in wire on right must be down direction

so along negative Y-direction

5)

force by the wire on left = force by wire on right

(\mu/4\pi) (2I1 I2 H/L) - (\mu/4\pi) (2I1 I2 H/(L+ W)) = (\mu/4\pi) (2I3 I2 H/(2L)) - (\mu/4\pi) (2I3 I2 H/(2L+ W))

(I1/L) - (I1/(L + W)) = (I3/(2L)) - (I3/(2L + W))

(0.575/0.20) - ((0.575)/(0.20 + 0.79)) = (I3/(0.40)) - (I3/(0.40 + 0.79))

I3 = 1.4 A